OpenStudy (anonymous):

You fill a container 3/4 full of water. The amount of water now in the container is 6 quarts. How much can the container hold?

OpenStudy (anonymous):

@DarkMoonZ

OpenStudy (bloomlocke367):

Okay, the easiest thing to do would to set up a proportion... do you know how to do that?

OpenStudy (anonymous):

um...no i don't

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (bloomlocke367):

Okay. well there are 2 different ones you could do:\[\frac{ .75 }{ 1 }=\frac{ 6 }{? }\] or you could do \[\frac{ 3 }{ 4 }=\frac{ 6 }{ ? }\] and get the same answer. You have to cross multiply and divide.

OpenStudy (anonymous):

ok

OpenStudy (bloomlocke367):

let me know what you get and I'll tell you if it's correct. Do you understand why I set them up like that?

OpenStudy (anonymous):

ok so is 8?

OpenStudy (anonymous):

and yes i understand now

OpenStudy (bloomlocke367):

yes :) good job.